# A stochastic method to estimate a zero-inflated two-part mixed model for human microbiome data

### v1

@@ 2.3 The SAEM algorithm for ZIBR parameter estimation
[…]
$\begin{split}p(\bm{y},\bm{\varphi};\theta)=&p(\bm{y}|\bm{\varphi};\alpha,\beta,\phi)p(\bm{\varphi}|\bm{\mu},\mathbf{G})\\ \propto&|\mathbf{G}|^{-\frac{N}{2}}\prod_{i}\exp{\left(-\frac{(\varphi_{i}-\bm{\mu})^{T}\mathbf{G^{-1}}(\varphi_{i}-\bm{\mu})}{2}\right)}\\ &\times\prod_{i,t}(1-p_{it})^{\mathbbm{1}_{\left\{y_{it}=0\right\}}}p_{it}^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}f(y_{it};u_{it},\phi)^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}.\end{split}$ (5)
Like most zero-inflated models, the ZIBR model cannot be considered part of the exponential family (Eggers,, 2015). However, the decomposition presented in Equation 5 allows us to propose a simplified structure for the SAEM algorithm (Equation 4). For the multivariate normal part corresponding to the random effects, the actualization in the SA step is done on the respective sufficient statistics. For the mixture distribution corresponding to the observed data, $\bm{y}|\bm{\varphi};\alpha,\beta,\phi$, maximization of the conditional log-likelihood is followed by estimates updates, as suggested for non-exponential family models (Comets et al.,, 2021). Then, the Maximum Likelihood iterative estimation algorithm for the parameters of the ZIBR model writes, for a given starting point $\theta^{(0)}$ and at iteration $q$, as:
@@
    Simulation step: draw $\varphi^{(q)}_{i},i=1,\cdots,N$ from the distribution $p(\cdot|\bm{y};\theta^{(q-1)})$.
@@
    Stochastic Approximation step: update the summary data functions $F_{1}(\bm{y},\bm{\varphi})$ and $F_{2}(\bm{y},\bm{\varphi})$ with the scheme:
    $\begin{split}F_{1}^{(q)}(\bm{y},\bm{\varphi})&=F_{1}^{(q-1)}(\bm{y},\bm{\varphi})+\gamma_{q}\left(\sum_{i}{\varphi_{i}^{(q)}}-F_{1}^{(q-1)}(\bm{y},\bm{\varphi})\right)\\ F_{2}^{(q)}(\bm{y},\bm{\varphi})&=F_{2}^{(q-1)}(\bm{y},\bm{\varphi})+\gamma_{q}\left(\sum_{i}{\varphi_{i}^{(q)}{\varphi_{i}^{(q)}}^{T}}-F_{2}^{(q-1)}(\bm{y},\bm{\varphi})\right).\end{split}$ (6)
@@
    $\begin{split}\bm{\mu}^{(q)}&=\frac{F_{1}^{(q)}(\bm{y},\bm{\varphi})}{N}\\ \mathbf{G}^{(q)}&=\frac{F_{2}^{(q)}(\bm{y},\bm{\varphi})}{N}-\frac{F_{1}^{(q)}(\bm{y},\bm{\varphi}){F_{1}^{(q)}(\bm{y},\bm{\varphi})}^{T}}{N^{2}}\end{split}$ (7)
@@
    $\left(\tilde{\beta}^{(q)},\tilde{\phi}^{(q)}\right)=$ (8)
    $\operatorname*{arg\,max}_{\beta,\phi}\sum_{i,t}$ $\left[\mathbbm{1}_{\left\{y_{it}>0\right\}}\left(\log{\frac{\Gamma(\phi)}{\Gamma\left(u_{it}^{(q)}\phi\right)\Gamma\left(\left(1-u_{it}{(q)}\right)\phi\right)}}+u_{it}^{(q)}\phi\log{y_{it}}+\phi\left(1-u_{it}^{(q)}\right)\log{(1-y_{it})}\right)\right]$
@@
    $\tilde{\alpha}^{(q)}=\operatorname*{arg\,max}_{\alpha}\sum_{i,t}\left[\mathbbm{1}_{\left\{y_{it}>0\right\}}\log{\left(p_{it}^{(q)}\right)}+\mathbbm{1}_{\left\{y_{it}=0\right\}}\log{\left(1-p_{it}^{(q)}\right)}\right].$ (9)
    where $u_{it}^{(q)}=u_{it}^{(q)}(b_{i},\beta)$ and $p_{it}^{(q)}=p_{it}^{(q)}(a_{i},\alpha)$ are calculated using $\varphi^{(q)}_{i}$ and Equation 2. Maximization in (8) and (9) is achieved numerically. Finally, the values are updated by doing
    $\begin{split}\phi^{(q)}&=\phi^{(q-1)}+\gamma_{q}\left(\tilde{\phi}^{(q)}-\phi^{(q-1)}\right)\\ \alpha^{(q)}&=\alpha^{(q-1)}+\gamma_{q}\left(\tilde{\alpha}^{(q)}-\alpha^{(q-1)}\right)\\ \beta^{(q)}&=\beta^{(q-1)}+\gamma_{q}\left(\tilde{\beta}^{(q)}-\beta^{(q-1)}\right)\end{split}$ (10)
Let us discuss the details of this implementation. As mentioned in 2.2, the choice of the starting point $\theta^{(0)}$ for SAEM does not affect its convergence; however, it is recommended to use values obtained in previous studies or with other estimation methods. Following the example of the existing implementation of the saemix packageComets et al., (2017), we will use $\gamma_{q}$ defined as follows:
$\gamma_{q}=\begin{cases}1&\mbox{if $q\leq K_{1}$,}\\ \frac{1}{q-K_{1}}&\mbox{if $K_{1}<q\leq K_{1}+K_{2}$.}\\ \end{cases}$
where $K_{1}+K_{2}$ is the total number of iterations.
We also discussed in section 2.2 that it is possible to improve the performance of the algorithm by taking multiple sequences or Markov chains in the Simulation step, and using Monte Carlo in Equations 7 and 10. Furthermore, during the SA step, we obtain sequences that allow to estimate $\mathbb{E}\left(\varphi_{i}|y_{i};\hat{\theta}\right)$ and $\mbox{Var}\left(\varphi_{i}|y_{i};\hat{\theta}\right)$ to be calculated, values necessary to approximate the log-likelihood through Importance Sampling, with which the Likelihood Ratio Test (LRT) can be computed, as presented in the following subsection.
@@ 2.3.1 Approximation of the log-likelihood using Importance Sampling
The log-likelihood of the observed data cannot be computed in closed form for complex mixed effects models, but its estimation is required to perform the LRT and to compute information criteria for a given model. One approximation method is given by the application of the Importance Sampling algorithm (Kloek and van Dijk,, 1978). Let $\mathcal{L}\mathcal{L}_{y}(\hat{\theta})$ be the log-likelihood of the model at the vector of population parameter estimates, that is $\mathcal{LL}_{y}(\hat{\theta})=\log p(\bm{y};\hat{\theta})$ where $p(\bm{y};\hat{\theta})=L(\hat{\theta};\bm{y})$ is the joint probability distribution function of the observed data given $\hat{\theta}$. Notice that $\mathcal{LL}_{y}(\hat{\theta})=\log p(\bm{y};\hat{\theta})=\sum_{i=1}^{N}\log p(y_{i};\hat{\theta})$ and, for some proposal distribution $\tilde{p}_{\varphi_{i}}$ absolutely continuous with respect to $p_{\varphi_{i}}$, we have
@@
That is, $p(y_{i};\hat{\theta})$ can be expressed as an expectation which can be approximated by:
@@
    Obtain a random sample of size $K$ $\varphi_{i}^{(1)},\cdots,\varphi_{i}^{(K)}$ from the proposal distribution $\tilde{p}_{\varphi_{i}}$;
@@
    Compute the empirical mean $\hat{p}_{(i,K)}=\frac{1}{K}\sum_{k=1}^{K}p(y_{i}|\varphi_{i}^{(k)};\hat{\theta})\frac{p(\varphi_{i}^{(k)};\hat{\theta})}{\tilde{p}(\varphi_{i}^{(k)};\hat{\theta})}$
An optimal proposal distribution would be the conditional distribution $p_{\varphi_{i}|y_{i}}$ since in that case the estimator of the expectation has zero variance. But since the closed form expression of the distribution is not available, we choose a proposal close to this optimal distribution, based on the empirically estimated conditional mean and variance, $\mu_{i}=\hat{\mathbb{E}}\left[\varphi_{i}|y_{i};\hat{\theta}\right]$ and $\sigma^{2}_{i}=\hat{\mbox{Var}}\left[\varphi_{i}|y_{i};\hat{\theta}\right]$, of the simulated random effects during the simulation step of the SAEM algorithm. Then, the proposed candidate $\varphi_{i}^{(k)}$, with $k=1,\dots,K$, is drawn with a noncentral student $t$-distribution $\varphi_{i}^{(k)}=\mu_{i}+\sigma_{i}\times T_{i,k}$, with $T_{i,k}\sim t_{\nu}$ i.i.d., where $t_{\nu}$ denotes a Student’s t-distribution with $\nu$ degrees of freedom. In this work, unless otherwise mentioned, the parameters for calculating the log-likelihood will be $\nu=5$ and $K=500$.
@@ 2.3.2 Stochastic approximation of the standard errors
$\partial_{\theta}^{2}\log p(\bm{y};\theta)=\mathbb{E}\left(\partial_{\theta}^{2}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)+\mbox{Cov}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)$
@@
$\mbox{Cov}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)=$ $\mathbb{E}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)^{T}|\bm{y};\theta\right)$
$-\mathbb{E}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)\mathbb{E}\left(\partial_{\theta}\log p(\bm{y},\bm{\varphi};\theta)|\bm{y};\theta\right)^{T}$
Given this, the second order derivative of the observed likelihood function with respect to parameter $\hat{\theta}$, $\partial_{\theta}^{2}L(\hat{\theta};\bm{y})$, can be approximated by the sequence $\{H_{q}\}_{q\in\mathbb{N}}$ which is calculated at iteration $q$ of the SAEM algorithm as:
@@
$H_{q}$ $=$ $G_{q}-D_{q}D_{q}^{\prime}.$
At convergence, $-H_{q}^{-1}$ can be used to approximate the covariance matrix of the parameter estimates (Zhu and Lee,, 2002; Cai,, 2010), which are useful to derivate procedures for hypothesis testing for the different parameters of the model.
@@ 3 Simulation studies
To evaluate the behavior of the proposed estimation method, and to compare it with existing alternatives, we conducted several simulation studies. It is worth noticing that the GHQ approach does not allow to deal with a different number of observations per individual, which is possible with our SAEM-based approach and the gamlss package. Therefore, we present simulations with balanced data first. Additional simulations on unbalanced data are also provided in Appendix A, Supplementary Materials, in which the performance of SAEM on the unbalanced datasets is compared with the use of the GHQ algorithm on balanced datasets obtained from imputation, and with gamlss without imputation. Covariates significance analysis based on the LRT and the Wald test are also presented in Appendix B, Supplementary Materials.
@@ 3.1 Setup
  • Setting 1: $a=b=-0.5$, $\alpha=\beta=0.5$, $\sigma_{1}=3.2$, $\sigma_{2}=2.6$, $\phi=6.4$.
  • Setting 2: $a=b=-0.5$, $\alpha=\beta=0.5$, $\sigma_{1}=0.7$, $\sigma_{2}=0.5$, $\phi=6.4$.
@@
For both Settings 1 and 2, $R=1000$ datasets were generated, and the SAEM estimation was implemented with $m=5$ chains and $(K_{1},K_{2})=(750,250)$, having therefore 1000 total iterations, with a starting point $\theta_{0}=(\phi_{0},a_{0},b_{0},\alpha_{0},\beta_{0},\sigma_{1,0},\sigma_{2,0})=(8,-0.3,-0.2,0.7,0.8,0.38,0.31)$.

### v2

@@ 2.3 The SAEM algorithm for ZIBR parameter estimation
As we have seen before, a mixed model can be considered as an unobserved data problem and therefore be addressed using the SAEM algorithm. Let us consider $\bm{\varphi}_{i}=(a_{i},b_{i})$, $1\leq i\leq N$, the non-observed data. By the definition of the ZIBR model $\bm{\varphi}_{i}$ follows the multivariate normal distribution $\bm{\varphi}_{i}\sim N(\bm{\mu},\mathbf{G})$ with $\bm{\mu}=(a,b)$ and $\mathbf{G}=diag(\sigma_{1}^{2},\sigma_{2}^{2})$. With the usual notation $\bm{y}=(y_{it}:\ 1\leq i\leq N,1\leq t\leq T_{i})$ and $\bm{\varphi}=(\bm{\varphi}_{i}:\ 1\leq i\leq N)$, the complete-data likelihood writes:
$\begin{split}p(\bm{y},\bm{\varphi};\theta)=&p(\bm{y}|\bm{\varphi};\alpha,\beta,\phi)p(\bm{\varphi}|\bm{\mu},\mathbf{G})\\ \propto&|\mathbf{G}|^{-\frac{N}{2}}\prod_{i}\exp{\left(-\frac{(\bm{\varphi}_{i}-\bm{\mu})^{T}\mathbf{G^{-1}}(\bm{\varphi}_{i}-\bm{\mu})}{2}\right)}\\ &\times\prod_{i,t}(1-p_{it})^{\mathbbm{1}_{\left\{y_{it}=0\right\}}}p_{it}^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}f(y_{it};u_{it},\phi)^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}.\end{split}$ (5)
Like most zero-inflated models, the ZIBR model cannot be considered part of the exponential family (Eggers,, 2015). However, the decomposition presented in Equation 5 allows us to propose a simplified structure for the SAEM algorithm (Equation 4). For the multivariate normal part corresponding to the random effects, the actualization in the SA step is done on the respective sufficient statistics. For the mixture distribution corresponding to the observed data, $\bm{y}|\bm{\varphi};\bm{\alpha},\bm{\beta},\phi$, maximization of the conditional log-likelihood is followed by estimates updates, as suggested for non-exponential family models (Comets et al.,, 2021). Then, the Maximum Likelihood iterative estimation algorithm for the parameters of the ZIBR model writes, for a given starting point $\bm{\theta}^{(0)}$, the summary data functions $F_{1}^{(0)}(\bm{y},\bm{\varphi})=\sum_{i}{\bm{\varphi}_{i}^{(0)}}$ and $F_{2}^{(0)}(\bm{y},\bm{\varphi})=\sum_{i}{\bm{\varphi}_{i}^{(0)}{\bm{\varphi}_{i}^{(0)}}^{T}}$ and at iteration $q$, as:
@@
    Simulation step: draw $\bm{\varphi}^{(q)}_{i},i=1,\cdots,N$ from the distribution $p(\cdot|\bm{y};\bm{\theta}^{(q-1)})$.
@@
    Stochastic Approximation step: update the summary data functions with the scheme:
    $\begin{split}F_{1}^{(q)}(\bm{y},\bm{\varphi})&=F_{1}^{(q-1)}(\bm{y},\bm{\varphi})+\gamma_{q}\left(\sum_{i}{\bm{\varphi}_{i}^{(q)}}-F_{1}^{(q-1)}(\bm{y},\bm{\varphi})\right)\\ F_{2}^{(q)}(\bm{y},\bm{\varphi})&=F_{2}^{(q-1)}(\bm{y},\bm{\varphi})+\gamma_{q}\left(\sum_{i}{\bm{\varphi}_{i}^{(q)}{\bm{\varphi}_{i}^{(q)}}^{T}}-F_{2}^{(q-1)}(\bm{y},\bm{\varphi})\right).\end{split}$ (6)
@@
    $\begin{split}\bm{\mu}^{(q)}&=\frac{F_{1}^{(q)}(\bm{y},\bm{\varphi})}{N}\\ \mathbf{G}^{(q)}&=\frac{F_{2}^{(q)}(\bm{y},\bm{\varphi})}{N}-\frac{F_{1}^{(q)}(\bm{y},\bm{\varphi}){F_{1}^{(q)}(\bm{y},\bm{\varphi})}^{T}}{N^{2}}.\end{split}$ (7)
@@
$\left(\tilde{\bm{\beta}}^{(q)},\tilde{\phi}^{(q)}\right)=$ (8)
$\operatorname*{arg\,max}_{\bm{\beta},\phi}\sum_{i,t}$ $\left[\mathbbm{1}_{\left\{y_{it}>0\right\}}\left(\log{\frac{\Gamma(\phi)}{\Gamma\left(u_{it}^{(q)}\phi\right)\Gamma\left(\left(1-u_{it}^{(q)}\right)\phi\right)}}+u_{it}^{(q)}\phi\log{y_{it}}+\phi\left(1-u_{it}^{(q)}\right)\log{(1-y_{it})}\right)\right]$
@@
$\tilde{\bm{\alpha}}^{(q)}=\operatorname*{arg\,max}_{\bm{\alpha}}\sum_{i,t}\left[\mathbbm{1}_{\left\{y_{it}>0\right\}}\log{\left(p_{it}^{(q)}\right)}+\mathbbm{1}_{\left\{y_{it}=0\right\}}\log{\left(1-p_{it}^{(q)}\right)}\right]$ (9)
where $u_{it}^{(q)}=u_{it}^{(q)}(b_{i},\bm{\beta})$ and $p_{it}^{(q)}=p_{it}^{(q)}(a_{i},\bm{\alpha})$ are calculated using $\varphi^{(q)}_{i}$ and Equation 2. Maximization in (8) and (9) is achieved numerically. Finally, the values are updated by doing
$\begin{split}\phi^{(q)}&=\phi^{(q-1)}+\gamma_{q}\left(\tilde{\phi}^{(q)}-\phi^{(q-1)}\right)\\ \bm{\alpha}^{(q)}&=\bm{\alpha}^{(q-1)}+\gamma_{q}\left(\tilde{\bm{\alpha}}^{(q)}-\bm{\alpha}^{(q-1)}\right);\quad\bm{\beta}^{(q)}=\bm{\beta}^{(q-1)}+\gamma_{q}\left(\tilde{\bm{\beta}}^{(q)}-\bm{\beta}^{(q-1)}\right)\end{split}$ (10)
Let us discuss the details of this implementation. As mentioned in 2.2, the choice of the starting point $\bm{\theta}^{(0)}$ for SAEM does not affect its convergence; however, it is recommended to use values obtained in previous studies or with other estimation methods. Following the example of the existing implementation of the saemix package (Comets et al.,, 2017), we will use $\gamma_{q}=1$ if $q\leq K_{1}$ , and $\gamma_{q}=\frac{1}{q-K_{1}}$ if $K_{1}<q\leq K_{1}+K_{2}$, where $K_{1}+K_{2}$ is the total number of iterations.
[…]
@@ 2.3.1 Approximation of the log-likelihood using Importance Sampling
$p(y_{i};\hat{\bm{\theta}})=\int p(y_{i},\bm{\varphi}_{i};\hat{\bm{\theta}})d\bm{\varphi}_{i}=\int p(y_{i}|\bm{\varphi}_{i};\hat{\bm{\theta}})\frac{p(\bm{\varphi}_{i};\hat{\bm{\theta}})}{\tilde{p}(\bm{\varphi}_{i};\hat{\bm{\theta}})}\tilde{p}(\bm{\varphi}_{i};\hat{\bm{\theta}})d\bm{\varphi}_{i}=\mathbb{E}_{\tilde{p}}\left[p(y_{i}|\bm{\varphi}_{i};\hat{\bm{\theta}})\frac{p(\bm{\varphi}_{i};\hat{\bm{\theta}})}{\tilde{p}(\bm{\varphi}_{i};\hat{\bm{\theta}})}\right$].
[…]
@@
    Obtain a random sample $\bm{\varphi}_{i}^{(1)},\cdots,\bm{\varphi}_{i}^{(K)}$ from the proposal distribution $\tilde{p}_{\bm{\varphi}_{i}}$;
@@
    Compute the empirical mean $\hat{p}_{(i,K)}=\frac{1}{K}\sum_{k=1}^{K}p(y_{i}|\bm{\varphi}_{i}^{(k)};\hat{\bm{\theta}})\frac{p(\bm{\varphi}_{i}^{(k)};\hat{\bm{\theta}})}{\tilde{p}(\bm{\varphi}_{i}^{(k)};\hat{\bm{\theta}})}$
An optimal proposal distribution would be the conditional distribution $p_{\bm{\varphi}_{i}|y_{i}}$ since in that case the estimator of the expectation has zero variance. But since the closed form expression of the distribution is not available, we choose a proposal close to this optimal distribution, based on the empirically estimated conditional mean and variance, $\mu_{i}=\hat{\mathbb{E}}\left[\bm{\varphi}_{i}|y_{i};\hat{\bm{\theta}}\right$] and $\sigma^{2}_{i}=\hat{\mbox{Var}}\left[\bm{\varphi}_{i}|y_{i};\hat{\bm{\theta}}\right$], of the simulated random effects during the simulation step of the SAEM algorithm. Then, the proposed candidate $\bm{\varphi}_{i}^{(k)}$, with $k=1,\dots,K$, is drawn with a noncentral Student $t$-distribution $\bm{\varphi}_{i}^{(k)}=\mu_{i}+\sigma_{i}\times T_{i,k}$, with $T_{i,k}\sim t_{\nu}$ i.i.d., where $t_{\nu}$ denotes a Student t-distribution with $\nu$ degrees of freedom. In this work, unless otherwise mentioned, the parameters for calculating the log-likelihood will be $\nu=5$ and $K=500$.
@@ 2.3.2 Stochastic approximation of the standard errors
$\partial_{\theta}^{2}\log p(\bm{y};\bm{\theta})=\mathbb{E}\left(\partial_{\bm{\theta}}^{2}\log p(\bm{y},\bm{\varphi};\bm{\theta})|\bm{y};\bm{\theta}\right)+\mbox{Cov}\left(\partial_{\bm{\theta}}\log p(\bm{y},\bm{\varphi};\bm{\theta})|\bm{y};\bm{\theta}\right)$
@@
$\mbox{Cov}\left(\partial_{\bm{\theta}}\log p(\bm{y},\bm{\varphi};\bm{\theta})|\bm{y};\bm{\theta}\right)=$ $\mathbb{E}\left(\partial_{\bm{\theta}}\log p(\bm{y},\bm{\varphi};\bm{\theta})\partial_{\bm{\theta}}\log p(\bm{y},\bm{\varphi};\bm{\theta})^{\top}|\bm{y};\bm{\theta}\right)$
$-\mathbb{E}\left(\partial_{\bm{\theta}}\log p(\bm{y},\bm{\varphi};\bm{\theta})|\bm{y};\bm{\theta}\right)\mathbb{E}\left(\partial_{\bm{\theta}}\log p(\bm{y},\bm{\varphi};\bm{\theta})|\bm{y};\bm{\theta}\right)^{\top}$
Given this, the second order derivative of the observed likelihood function with respect to parameter $\hat{\bm{\theta}}$, $\partial_{\bm{\theta}}^{2}L(\hat{\bm{\theta}};\bm{y})$, can be approximated by the sequence $\{H_{q}\}_{q\in\mathbb{N}}$ which is calculated at iteration $q$ of the SAEM algorithm as:
@@
and $H_{q}=G_{q}-D_{q}D_{q}^{\prime}$. At convergence, $H_{q}^{-1}$ can be used to approximate the covariance matrix of the parameter estimates (Zhu and Lee,, 2002; Cai,, 2010), which are useful to derivate procedures for hypothesis testing for the different parameters of the model.
@@ 3 Simulation studies
To evaluate the behavior of the proposed estimation method, and to compare it with existing alternatives, we conducted several simulation studies. It is worth noticing that the GHQ approach does not allow to deal with a different number of observations per individual, which is possible with our SAEM-based approach and the gamlss package. Therefore, we present simulations with balanced data first. Additional simulations on unbalanced data are also provided in Appendix A, in which the performance of SAEM on the unbalanced datasets is compared with the use of the GHQ algorithm on balanced datasets obtained from imputation, and with gamlss without imputation. Covariates significance analysis based on the LRT and the Wald test are also presented in Appendix B. To evaluate how the statistical power of the SAEM and GHQ methods varies with the proportion of zeros in the data, we performed a set of simulation experiments. The corresponding results are reported in Appendix C.
@@ 3.1 Setup
  • Setting 1: a=b=0.5a=b=-0.5, $\alpha=\beta=0.5$, $\sigma_{1}=3.2$, $\sigma_{2}=2.6$, $\phi=6.4$.
  • Setting 2: a=b=0.5a=b=-0.5, $\alpha=\beta=0.5$, $\sigma_{1}=0.7$, $\sigma_{2}=0.5$, $\phi=6.4$.
@@
For both Settings 1 and 2, $R=1000$ datasets were generated, and the SAEM estimation was implemented with $m=5$ chains and $(K_{1},K_{2})=(750,250)$, having therefore 1000 total iterations, with a starting point $\bm{\theta}_{0}=(\phi_{0},a_{0},b_{0},\alpha_{0},\beta_{0},\sigma_{1,0},\sigma_{2,0})=(8,-0.3,-0.2,0.7,0.8,0.38,0.31)$. These quantities were determined using a graphical approach based on the trace plots of the parameter estimates across iterations. Convergence was assessed through visual inspection of the parameter trajectories. Initially, a large value of $K_{1}$ (with $\gamma_{k}$=1 for all iterations up to $K_{1}$) was used. From these plots, an appropriate value of $K_{1}$ was then selected once the iterates began to fluctuate randomly around a stable region.